Twelve different video games showing substance use were observed and the duration times of game play​ (in seconds) are listed below. The design of the study justifies the assumption that the sample can be treated as a simple random sample. Use the data to construct a

9595​%

confidence interval estimate of

muμ​,

the mean duration of game play.

40464046

38783878

38523852

40174017

43254325

48154815

46514651

40354035

50135013

48174817

43284328

43194319

Click here to view a t distribution table.

LOADING...
Answer #1

Solution :

Given that 4046,3878,3852,4017,4325,4815,4651,4035,5013,4817,4328,4319

=> mean x-bar = sum of terms/number of terms

= 52096/12

= 4341.3333

=> standard deviation s = 397.6031

=> n = 12 , df = n - 1 = 11

=> For 95% confidence interval , t = 2.179

=> The 95% confidence interval of the mean is

=> x-bar +/- t*s/sqrt(n)

=> 4341.3333 +/- 2.179*397.6031/sqrt(12)

=> (4091.2318 , 4591.4348)